OpenStudy (anonymous):

What is the distance between points (5, 5) and (1, 5) on a coordinate plane?

OpenStudy (alexandervonhumboldt2):

use distance formula: sqrt((x1-x2)^2+(y1-y2)^2)

OpenStudy (alexandervonhumboldt2):

substitute your given points into formula and solve

OpenStudy (anonymous):

i have no clue

OpenStudy (alexandervonhumboldt2):

look: sqrt((x1-x2)^2+(y1-y2)^2)=sqrt((5-1)^2+(5-5)^2)=sqrt(16+0)=sqrt16=4 thus the distance is 4

OpenStudy (anonymous):

ty